OpenStudy (anonymous):

Integrating ky(1-y), where k is an unknown constant and the integral = to 1 between infinity and negative infinity

OpenStudy (anonymous):

k should be 6, but I am unsure about how exactly that comes to be

OpenStudy (experimentx):

what is between infinity and negative infinite ??

OpenStudy (anonymous):

This is a density function for a continuous stochastic variable, f(y) = ky(1-y), 0 <= y <= 1 I need to find k, for this function to be a probability density function

OpenStudy (experimentx):

sorry ... I have no experience with stochastic calculus. i thought it was elementary calculus question.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The stochastic part I can handle, it is the calculus that is confusing me. It should integrate to 1 in the end, and I just need the procedure to integrate that bad boy

OpenStudy (experimentx):

If it's just improper integration then maybe i could help could you type your equation on here http://www.codecogs.com/latex/eqneditor.php and post/paste it ... or use equation tool to type equation?

OpenStudy (anonymous):

http://latex.codecogs.com/gif.latex? \int%20ky(1-y)dy%20=%201

OpenStudy (experimentx):

is this the equation? \[ \int_0^\infty ky(1-y) dy\]

OpenStudy (anonymous):

\[\int\limits_{ - \infty}^{\infty}ky(1-y) dy = 1\] Excuse my ignorance, I'm used to pencil and paper

OpenStudy (experimentx):

this integral does not converge I don't think it has Cauchy Principle Value either.

OpenStudy (anonymous):

I know that k = 6, for the function to integrate to 1. But thank you anyway. I will go scribble on a wall somewhere to figure it out.

OpenStudy (experimentx):

sure ... good luck.

OpenStudy (anonymous):

Um, looking at the supposed answer, I think the integral should be \[\int_0^1 ky(1-y) dy\] Since \(0 \leq y \leq 1\) Then it would be \[ k * \left(\frac{y}{2} - \frac{y^3}{3}\right)\bigg|^1_0=\frac{k}{6} \] Then k would be 6 for the integral to evaluate to 1..